# Neural Network-Based Bandit: A Medium Access Control for the IIoT Alarm Scenario

### v1

@@ Abstract
Efficient Random Access (RA) is critical for enabling reliable communication in Industrial Internet of Things (IIoT) networks. Herein, we propose a deep reinforcement learning based distributed RA scheme, entitled Neural Network-Based Bandit (NNBB), for the IIoT alarm scenario. In such a scenario, the devices may detect a common critical event, and the goal is to ensure the alarm information is delivered successfully from at least one device. The proposed NNBB scheme is implemented at each device, where it trains itself online and establishes implicit inter-device coordination to achieve the common goal. Devices can transmit simultaneously on multiple orthogonal channels and each possible transmission pattern constitutes a possible action for the NNBB, which uses a deep neural network to determine the action. Our simulation results show that as the number of devices in the network increases, so does the performance gain of the NNBB compared to the Multi-Armed Bandit (MAB) RA benchmark. For instance, NNBB experiences a 7% success rate drop when there are four channels and the number of devices increases from 10 to 60, while MAB faces a 25% drop.
@@ I INTRODUCTION
The Industrial Internet of Things (IIoT) incorporates intelligence and autonomy into manufacturing through dependable wireless connectivity, production line insights, smart alerts, and predictive maintenance via data acquisition and processing from numerous devices. Notably, the number of IIoT devices is expected to reach around $9.4$ billion by $2030$, representing $37\%$ of the total population of devices [1], and will lead to massive IIoT networks.
Some critical challenges in massive IIoT networks include wireless channel overload, packet loss due to high latency and deadline violation, and energy consumption escalation at the devices, e.g., due to repeated channel access requests. These challenges constitute the so-called massive access problem [2, 3]. In general, channel access solutions addressing this problem in IIoT deployments must adapt to sporadic and correlated traffic [4, 5], require low signaling [6, 5], support decentralized decision-making [7, 5], and be scalable [7, 5]. Orthogonal Multiple Access (MA) schemes cannot be used since they require excessive signaling, e.g., long orthogonal pilot sequences, and do not support decentralized decision-making. Although Aloha-based Random Access (RA) schemes satisfy the first three conditions, they are plagued by limited scalability, making them unfeasible in a massive IIoT network. This, and the boom of tinyML for on-device intelligence [8, 9], has motivated the research on learning-based RA for massive IIoT networks.
@@ I-A STATE-OF-ART ON LEARNING-BASED RA
The authors in [10] proposed a RL-based distributed RA scheme for Delay-Constrained (DC) communications (RLRA-DC). It establishes cooperation among devices to increase the system throughput beyond the maximum system throughput of $1/e$ achieved by the Aloha-based schemes. However, RLRA-DC requires information about the total number of devices present in the network. Although there are approaches to estimate such information, e.g., [10, 20], they are by no means error-free, inevitably leading to performance degradation of RLRA-DC in practice. The authors in [11] proposed a Tiny State-space R-learning random Access (TSRA) scheme to convey DC traffic. TSRA is a distributed RA scheme exploiting information regarding the urgency level of the data packets, which is measured in terms of their time to expire. The most important feature of TSRA is the independence of the size of the device state space from the number of devices in the network and the hard deadline for transmitting a packet. In spite of this, RLRA-DC achieves a higher system throughput than TSRA.
In [12], the authors considered a multi-access point (AP) scenario, where each user runs a two-stage RL based RA algorithm for maximizing the throughput. In the first stage, the user selects the AP at the beginning of the macro-time slot, while in the second stage, the user decides the sub-time slot for transmission. Despite its appealing distributed nature, the proposal has the following limitations: (i) a large number of devices in the network would make the lookup table of the second RL algorithm too large, thus taking a significant time to converge and degrading the solution’s optimality; and (ii) information about the devices associated with an AP is required after every macro-time slot.
The authors in [13] proposed a coordinated RA scheme adapted to sporadic and correlated IIoT traffic. Therein, the time slot selection task at each device is modelled as a Markov game, where devices are the agents, the number of retransmissions is the context, and the transmit time slots in a frame are the actions. Moreover, each device runs an RL-based Linear Reward-Inaction (LRI) algorithm [21] to learn the equilibrium points of the Markov game. Despite its appealing distributed nature, LRI in [13] never converged to a pure Nash equilibrium, leading to the coordinated RA scheme being beneficial only for moderate traffic conditions. Meanwhile, the authors in [14] developed a type of contention-cum-RL-based RA scheme called Inner-State-Driven random Access (ISDA). ISDA uses a Neural Network (NN) with one hidden layer of five neurons to determine the transmission probabilities of the devices in a time slot as a function of their inner states. The main benefits of ISDA include the decentralized modus operandi and the ability to meet heterogeneous performance requirements. However, its scalability is unproven as it was tested for a network with only three devices.
@@
Ref. Learning method Scenario
[13] LRI Uplink transmission of correlated IIoT traffic through a cellular network
[6] MAB Transmission of sensed alarm information to the ExC
[7] DRL Uplink transmission in a massive IoT network
[10, 11] R-learning Uplink transmission of delay-constrained traffic
@@
[12] MAB Uplink transmission in a multicell system
@@
[14] Cross-entropy [22] Network having a diverse quality of service demands from multiple IoT applications
[15] PPO Uplink transmission in heterogeneous IoT network
@@
[16] DQN Coexisting networks, each operating with a distinct medium access protocol
@@
[17] DQN Uplink transmission of sporadic traffic on a single channel in a slotted system
@@
[18] DQN Millimeter-wave network with numerous small/macro cells, consisting of traversing users
@@
[19] DQN Efficient spectrum access in a multichannel network with a simple collision model
@@
The authors in [7] proposed a multi-agent DRL-based RA scheme that employs a centralized training and decentralized execution approach for the uplink transmission in massive IoT setups. The learned RA policy in [7] is scalable, it can be shared among all the devices, and is adaptable to both correlated and regular IoT traffic. However, it has two flaws: (i) high complexity as devices must possess a NN with $256$ neurons in each of its two hidden layers, and (ii) convergence issues. To enhance the overall network throughput of a heterogeneous IoT network, the authors in [15] proposed a distributed Proximal Policy Optimization (PPO)-based MA (PPOMA). IoT nodes adopting PPOMA perform online learning to update their access policies and dynamically coexist with other nodes. A notable advantage of PPOMA is that it operates without demanding information about the MA protocols employed by the other coexisting IoT nodes. However, PPOMA requires three distinct NNs, namely actor, critic, and actor-target NNs, to be stored in the IoT node’s memory. Additionally, [15] acknowledges that while PPOMA enhances the overall network throughput, it does so at the expense of fairness among the coexisting IoT nodes.
@@
As mentioned earlier, DQN has also been adopted by many works. In [16], the authors introduced a DRL MA (DLMA) scheme exploiting DQN and aiming to maximize the sum throughput and the $\alpha$-fairness among the coexisting networks. DLMA allows the devices to learn how to properly use the time-spectral resources with no information about the medium access protocols used by other coexisting networks. However, DLMA demands an extensive NN with six hidden layers and $64$ neurons per each, and still cannot outperform the R-learning-based TSRA in terms of throughput, as stated in [11]. To learn a transmission policy that balances out throughput with fairness among users, the authors in [17] proposed a multi-agent DQN RA scheme. Here, fairness is measured via the age of a packet. The scheme employs a parameter-sharing method from [23] to train just one NN in a centralized manner and extend it to work for all devices. Unfortunately, even though the DQN RA scheme in [17] can tune itself to different Poisson-distributed data arrival rates, its performance degrades as the number of devices increases, as stated in [24].
@@
The authors in [18] proposed a DQN-based algorithm for wandering users to estimate the congestion levels of APs and select them accordingly, thus reducing the delay experienced by the users during contention-based RA. Meanwhile, the total number of successfully delivered packets was maximized in [19] by solving a dynamic spectrum access problem in a multichannel network. Because of the availability of many possible network states and the partial observability of those states due to zero message exchange among the devices, a pre-trained DQN-based RA scheme is proposed. The channel selection decisions are taken online and in a distributed fashion by the devices. However, a serious limitation is that the DQN at each device needs to be retrained offline every week or every time a significant environmental change occurs. It is worth noting that the DQN-based works [16, 18, 17, 19] have one limitation in common, which is that they all require a replica of the former NN model (also called target model) for training the latest NN model. Specifically, the approaches in [16, 18] require a replica to be stored in the device’s memory, while those in [17, 19] require it to be stored at the central unit.
@@ I-B MOTIVATION AND CONTRIBUTION
As discussed above, there are numerous learning-based RA schemes in the recent literature addressing several wireless communication scenarios. However, the IIoT alarm scenario remains much less explored. The alarm scenario consists of IoT devices that may sense the same critical event, such as an abnormal turbine vibration or an oil leak detection, and must promptly inform an External Controller (ExC). Hence, they must adopt a low-overhead protocol that avoids collisions from simultaneous alarm transmissions so that the critical alarm message transmission from at least one device succeeds.
@@
$Q(\boldsymbol{\mathring{a}}_{i})$ Action value of $\boldsymbol{\mathring{a}}_{i}$ in case of MAB
@@
In this work, we focus on the alarm scenario in an IIoT network. This problem has been studied earlier in [6], where the authors proposed a Multi-Armed Bandit (MAB) based RA scheme to allow the active devices, which detect the alarm, to indirectly coordinate their alarm transmissions to the ExC. The main limitations of the approach in [6] are: (i) no more than three devices can be active concurrently; (ii) training occurs at the active devices but only one at a time, which is time-consuming and imposes tight synchronization requirements on the network; and (iii) MAB does not perform well with a large action space. Meanwhile, existing Industrial Wireless Sensor Network (IWSN) protocols such as WirelessHART [25], ISA100.11100.11100.11100.11a [25], WIA-PA [25], employ Time Division MA (TDMA). TDMA is favored because it allows the prediction of communication latency for time-triggered packets in advance. However, centralized approaches like TDMA struggle to deliver event-triggered packets, such as IIoT alarm message packets, in a timely manner since the generation of the event-triggered packets is unpredictable [26]. The limitations of [6] and existing IWSN protocols motivate our work, which presents a novel, efficient, online learning-based RA for the alarm scenario. Our contributions are as follows:
@@
  • We devise a novel procedure to acquire a useful context to assist the learning-based RA. The procedure starts with the active devices transmitting their respective pilot signals to a Base Station (BS). These signals are received as an aggregated signal at the BS, which then broadcasts it to the active devices. At each device, this constitutes the context for the proposed RL framework.
@@
  • We propose a distributed learning-based RA scheme called Neural Network-Based Bandit (NNBB) for the IIoT alarm scenario. NNBB uses a simple Deep NN (DNN) to process the context received by a device and decide the transmission channel(s) for the alarm signals such that the ExC successfully receives the signal on at least one channel. Unlike [6], there is no restriction on the number of simultaneously active devices.
  • We devise an online training procedure for the NNBB in which the DNN weights are updated at every alarm event using Root Mean Square Propagation (RMSProp). Unlike DQN, NNBB does not require a target model for training.
@@
  • We compare NNBB with two benchmark schemes, namely MAB-based RA and Random Selection (RS), under various network configurations through simulations. Our results indicate that NNBB outperforms them even when the network experiences an increase in the available channels/devices or activation probability.

### v2

@@ Abstract
Efficient Random Access (RA) is critical for enabling reliable communication in Industrial Internet of Things (IIoT) networks. Herein, we propose a deep reinforcement learning-based distributed RA scheme, entitled Neural Network-Based Bandit (NNBB), for the IIoT alarm scenario. In such a scenario, devices may detect a common critical event, and the goal is to ensure the alarm information is delivered successfully from at least one device. The proposed NNBB scheme is implemented at each device, where it trains itself online and establishes implicit inter-device coordination to achieve the common goal. We devise a procedure for acquiring a valuable context for NNBB, which then uses a deep neural network to process this context and let devices determine their action. Each possible transmission pattern, i.e., transmit channel(s) allocation, constitutes a feasible action. Our simulation results show that as the number of devices in the network increases, so does the performance gain of the NNBB compared to the Multi-Armed Bandit (MAB) RA benchmark. For instance, NNBB experiences a 7% success rate drop when there are four channels and the number of devices increases from 10 to 60, while MAB faces a 25% drop.
@@ I INTRODUCTION
The Industrial Internet of Things (IIoT) incorporates intelligence and autonomy into manufacturing through dependable wireless connectivity, production line insights, smart alerts, and predictive maintenance via data acquisition and processing from numerous devices. Notably, the number of IIoT devices is expected to reach around $9.4$ billion by $2030$, representing $37\%$ of the total population of devices [1] and leading to massive IIoT networks.
Massive IIoT networks face several critical challenges, including wireless channel overload, packet loss caused by high latency and deadline violation, and increased energy consumption at devices, particularly due to repeated channel access requests. Collectively, these challenges constitute the so-called massive access problem [2, 3]. In general, channel access solutions addressing this problem in IIoT deployments must adapt to sporadic and correlated traffic [4, 5], require low signaling [6, 5], support decentralized decision-making [7, 5], and be scalable [7, 5]. Aloha-based Random Access (RA) schemes, such as pure Aloha and slotted Aloha, offer decentralized decision-making and are well-suited for sporadic and correlated traffic, with the added benefit of low signaling requirements. However, they suffer from limited scalability, especially in scenarios where IIoT devices become randomly active by sensing an alarm event, i.e., leading to highly heterogeneous activation profiles over time. This makes Aloha-based RA schemes impractical for massive IIoT networks, which together with the boom of tinyML for on-device intelligence [8, 9], has motivated the research on learning-based RA for massive IIoT networks. Learning-based RA schemes can effectively adapt to the spatial/temporal correlated nature of IIoT traffic, and even reduce the need for explicit signaling, as in [6]. Additionally, learning-based RA schemes can enable a decentralized modus operandi, which enhances the network’s ability to tolerate faults. In this decentralized modus operandi, each device independently learns its RA strategy, promoting scalability.
@@ I-A STATE-OF-ART ON LEARNING-BASED RA
Authors in [10] proposed an RL-based distributed RA scheme for Delay-Constrained (DC) communications (RLRA-DC). It establishes cooperation among devices to increase the system throughput beyond the maximum system throughput of $1/e$ achieved by the Aloha-based schemes. However, RLRA-DC requires information about the total number of devices in the network. Although there are approaches to estimate such information, e.g., [10, 20], they are by no means error-free, inevitably leading to performance degradation of RLRA-DC in practice. Authors in [11] proposed a Tiny State-space R-learning random Access (TSRA) scheme to convey DC traffic. TSRA is a distributed RA scheme exploiting information regarding the urgency level of the data packets, which is measured in terms of their time to expire. The most important feature of TSRA is the independence of the size of the device state space from the number of devices in the network and the hard deadline for transmitting a packet. In spite of this, RLRA-DC achieves a higher system throughput than TSRA.
In [12], authors considered a multi-access point (AP) scenario, where each user runs a two-stage RL based RA algorithm for maximizing the throughput. In the first stage, the user selects the AP at the beginning of the macro-time slot, while in the second stage, the user decides the sub-time slot for transmission. Despite its appealing distributed nature, the proposal has the following limitations: (i) a large number of devices in the network would make the lookup table of the second RL algorithm too large, thus taking a significant time to converge and degrading the solution’s optimality; and (ii) information about devices associated with an AP is required after every macro-time slot.
Authors in [13] proposed a coordinated RA scheme adapted to sporadic and correlated IIoT traffic. Therein, the time slot selection task at each device is modelled as a Markov game, where devices are the agents, the number of retransmissions is the context, and the transmit time slots in a frame are the actions. Moreover, each device runs an RL-based Linear Reward-Inaction (LRI) algorithm [21] to learn the equilibrium points of the Markov game. Despite its appealing distributed nature, LRI in [13] never converged to a pure Nash equilibrium, leading to the coordinated RA scheme being beneficial only for moderate traffic conditions. Meanwhile, authors in [14] developed a type of contention-cum-RL-based RA scheme called Inner-State-Driven random Access (ISDA). ISDA uses a Neural Network (NN) with one hidden layer of five neurons to determine the transmission probabilities of devices in a time slot as a function of their inner states. The main benefits of ISDA include the decentralized modus operandi and the ability to meet heterogeneous performance requirements. However, the following issue needs to be addressed [14]: the policy optimization method is highly sub-optimal.
@@
Ref. Learning method Scenario Limitations of corresponding RA schemes
[6] MAB Transmission of sensed alarm information to an external controller $\bullet$ No more than three devices can be active concurrently
$\bullet$ Training occurs at the active devices but only one at a time, which is time-consuming and imposes tight synchronization requirements on the network
$\bullet$ MAB does not perform well with a large action space
[7] DRL Uplink transmission in a massive IoT network $\bullet$ High complexity, as devices must possess a NN with $256$ neurons in each of its two hidden layers
$\bullet$ RA policy fails to converge when IoT devices lack information about other device’s identification
[10, 11] R-learning Uplink transmission of delay-constrained traffic $\bullet$ RA scheme in [10] requires information about the total number of devices present in the network
$\bullet$ Inaccuracies in the estimation of the total number of devices lead to the performance of degradation of RA scheme in [10]
$\bullet$ Even though RA scheme in [11] does not require information about the total number of devices, however, its system throughput is lower than the one obtained in [10]
@@
[12] MAB Uplink transmission in a multicell system $\bullet$ Takes significant time to converge
$\bullet$ Information about devices associated with an AP is required after every macro-time slot
[13] LRI Uplink transmission of correlated IIoT traffic through a cellular network $\bullet$ LRI never converges to a pure Nash equilibrium
$\bullet$ Performs badly for heavy traffic conditions
@@
[14] Cross-entropy [22] Network having a diverse quality of service demands from multiple IoT applications $\bullet$ The policy optimization method used in [14] is highly sub-optimal and it needs improvement
[15] PPO Uplink transmission in heterogeneous IoT network $\bullet$ Requires three distinct NNs to be stored in the IoT node’s memory
$\bullet$ Enhances the overall network throughput without any individual performance guarantees.
@@
[16] DQN Coexisting networks, each operating with a distinct medium access protocol $\bullet$ Demands an extensive NN with six hidden layers and $64$ neurons per each
$\bullet$ Throughput is lower than the one obtained in [11]
$\bullet$ Requires a replica of the former NN model (also called target model), to be stored in the device’s memory, for training the latest NN model
@@
[17] DQN Uplink transmission of sporadic traffic on a single channel in a slotted system $\bullet$ RA policy at each agent fails to converge due to a badly chosen reward function
$\bullet$ Requires a replica of the former NN model, to be stored at the central unit, for training the latest NN model
@@
[18] DQN Millimeter-wave network with numerous small/macro cells and traversing users $\bullet$ Requires a replica of the former NN model, to be stored in the device’s memory, for training the latest NN model
@@
[19] DQN Efficient spectrum access in a multichannel network with a simple collision model $\bullet$ DQN at each device needs to be retrained offline every time a significant environmental change occurs
$\bullet$ Requires a replica of the former NN model, to be stored at the central unit, for training the latest NN model
@@
Authors in [7] proposed a multi-agent DRL-based RA scheme that employs a centralized training and decentralized execution approach for uplink transmission in massive IoT setups. The learned RA policy in [7] is scalable, can be shared among all devices, and can adapt to both correlated and uncorrelated IoT traffic. However, the proposal has two flaws: (i) high complexity as devices must possess NN with $256$ neurons in each of its two hidden layers, and (ii) the RA policy fails to converge when IoT devices lack information about other device’s identification. To enhance the overall network throughput of a heterogeneous IoT network, authors in [15] proposed a distributed Proximal Policy Optimization (PPO)-based Multiple Access (MA). IoT nodes adopting PPOMA perform online learning to update their access policies and dynamically coexist with other nodes. A notable advantage of PPOMA is that it operates without demanding information about the MA protocols employed by the other coexisting IoT nodes. However, PPOMA requires three distinct NNs, namely actor, critic, and actor-target NNs, to be stored in the IoT node’s memory. Additionally, [15] acknowledges that while PPOMA enhances the overall network throughput, it does so without providing any individual performance guarantees.
@@
As mentioned earlier, DQN has also been adopted by many works. In [16], authors introduced a DRL MA (DLMA) scheme exploiting DQN and aiming to maximize the sum throughput and the $\alpha$-fairness among the coexisting networks. DLMA allows devices to learn how to properly use the time-spectral resources with no information about the medium access protocols used by other coexisting networks. However, DLMA demands an extensive NN with six hidden layers and $64$ neurons per each, and still cannot outperform the R-learning-based TSRA in terms of throughput, as stated in [11]. To learn a transmission policy that balances out throughput and fairness among users, authors in [17] proposed a multi-agent DQN RA scheme. In this work, fairness is impacted by the capture effect, which occurs when one user occupies the channel for an extended period, limiting other nodes’ transmission opportunities and causing longer delays. The capture effect is measured by the average packet age. The scheme employs a parameter-sharing method from [23] to train just one NN in a centralized manner and extends it to work for all devices. Unfortunately, even though the RA policy at an agent can tune itself to different Poisson-distributed data arrival rates, it fails to converge because of a badly chosen reward function, as stated in [24].
@@
Authors in [18] proposed a DQN-based algorithm for wandering users to estimate the congestion levels of APs and select them accordingly, thus reducing the delay experienced by the users during contention-based RA. Meanwhile, the total number of successfully delivered packets was maximized in [19] by solving a dynamic spectrum access problem in a multichannel network. Due to the availability of many possible network states and the partial observability of those states due to zero message exchange among devices, a pre-trained DQN-based RA scheme is proposed. The channel selection decisions are taken online and in a distributed fashion by devices. However, a serious limitation is that the DQN at each device needs to be retrained offline every week or every time a significant environmental change occurs. It is worth noting that the DQN-based works [16, 18, 17, 19] have one limitation in common: they all require a replica of the former NN model (also called target model) for training the latest NN model. Specifically, the approaches in [16, 18] require a replica to be stored in the device’s memory, while those in [17, 19] require it to be stored at the central unit.
@@ I-B MOTIVATION AND CONTRIBUTION
As discussed above, there are numerous learning-based RA schemes in the recent literature addressing several wireless communication scenarios. However, the IIoT alarm scenario remains much less explored. The alarm scenario consists of IoT devices that may sense the same critical event, such as a gas leak, temperature anomaly, or radiation leak, and must promptly inform an External Controller (ExC). Hence, they must adopt a low-overhead protocol that avoids collisions from simultaneous alarm transmissions so that the critical alarm message transmission from at least one device succeeds.
@@
$\vartheta_{lb},\vartheta_{ub}$ Lower and upper bound of NNBB’s complexity
@@
$\boldsymbol{\theta}_{\upsilon}$ Weight vector of active agent $\upsilon$ in the case of MQLFA
@@
$Q_{1}(\boldsymbol{\mathring{a}}_{i})$ Action value of $\boldsymbol{\mathring{a}}_{i}$ in the case of MAB
${Q_{2}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i},\boldsymbol{\theta}_{\upsilon})}$ Action value of $\boldsymbol{\mathring{a}}_{i}$, given $\boldsymbol{s}_{\upsilon}$ and $\boldsymbol{\theta}_{\upsilon}$, in the case of MQLFA
@@
$\boldsymbol{\phi}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i})$ Feature vector of MQLFA
@@
In this work, we focus on the alarm scenario in an IIoT network. This problem has been studied earlier in [6], where authors proposed a Multi-Armed Bandit (MAB) based RA scheme that allows devices detecting the alarm to indirectly coordinate their alarm transmissions to the ExC. The main limitations of the approach in [6] are: (i) it works under the assumption that no more than three devices can be active concurrently, while in practice it is not possible to control the number of active devices; (ii) training occurs at the active devices but only one at a time, which is time-consuming and imposes tight synchronization requirements on the network; and (iii) MAB does not perform well with a large action space. Meanwhile, existing Industrial Wireless Sensor Network (IWSN) protocols such as WirelessHART, ISA100.11100.11100.11100.11a, WIA-PA [25], employ Time Division MA (TDMA). TDMA is favored because it allows the prediction of communication latency for time-triggered packets in advance. However, centralized approaches like TDMA struggle to deliver event-triggered packets, such as IIoT alarm message packets, in a timely manner since the generation of the event-triggered packets is unpredictable [26]. Moreover, in an IIoT alarm scenario, a large number of devices may become active simultaneously depending on device density, thus, the corresponding RA scheme must account for this. The limitations of [6] and existing IWSN protocols in these regards motivate our work, which presents a novel, efficient, online learning-based RA for the alarm scenario. Our specific contributions are as follows:
@@
  • We devise a novel procedure to acquire a useful context to assist the learning-based RA. The procedure starts with the active devices transmitting their respective pilot signals to their Base Station (BS). These signals are received as an aggregated signal at the BS, which then broadcasts it to devices. At each active device, this constitutes the context for the proposed RL framework.
@@
  • We propose a distributed RL-based RA scheme for the IIoT alarm scenario, referred to as Neural Network-Based Bandit (NNBB). NNBB uses a simple Deep NN (DNN) to process the context received by a device and decide the transmission channel(s) for the alarm signals such that the ExC successfully receives the signal on at least one channel. Unlike [6], herein there is no restriction on the number of simultaneously active devices. Moreover, NNBB does not require information about the total number of devices in an IIoT network. Note that relevant limitations for on-device context processing to the potential alternatives of DNN, such as attention-based and diffusion models, as in [27, 28, 29], advise against their use here. Specifically, the attention mechanism aggravates the convergence, complexity, and computational cost of the hosting RL algorithm [30]. Meanwhile, diffusion models require numerous denoising steps during inference to generate a single action selection probability vector. This not only increases the computational complexity of diffusion models but also leads to significant overhead [29, 31].
  • We devise an online training procedure for NNBB in which the DNN weights are updated at every alarm event using Root Mean Square Propagation (RMSProp). Online training allows IIoT devices to adjust their actions almost immediately based on feedback from the ExC, which is crucial in industrial environments where delays can be costly. Additionally, the use of gradient clipping in NNBB’s training procedure facilitates the smooth convergence of our multi-agent distributed RL system.
@@
  • We compare NNBB with three benchmark schemes, namely MAB-based RA, Random Selection (RS), and Myopic Q-learning with Linear Function Approximation (MQLFA)-based RA, under various network configurations. Our simulation results indicate that NNBB outperforms them even when the network experiences an increase in the available channels/devices or activation probability.